OpenStudy (anonymous):

A company that packages cold cuts needs to transport the packaged meat to the supermarket. It will cost $1,200 to rent a truck, and it will cost an additional $100 for every 50 pounds transported. Which equation represents the transport cost for every 50 pounds of cold cuts transported?

OpenStudy (anonymous):

y = 1,200 − 100x y = 1,200 + 100x y = 1,200x + 100 y = 100x − 1,200

OpenStudy (anonymous):

ok what would it mean if it costed an additional 100 dollors, what is that saying

OpenStudy (anonymous):

You add 100

OpenStudy (anonymous):

which is the only one that says (+ 100)

OpenStudy (anonymous):

Huh??

OpenStudy (anonymous):

theres 2

OpenStudy (anonymous):

the one without the x is the one

OpenStudy (anonymous):

y = 1,200x + 100